### v1

# Negation: A Pink Elephant in the Large Language Models’ Room?

Abstract
Negations are key to determining sentence meaning, making them essential for logical reasoning. Despite their importance, negations pose a substantial challenge for large language models (LLMs) and remain underexplored.
We construct two multilingual natural language inference (NLI) datasets with paired examples differing in negation. We investigate how model size and language impact its ability to handle negation correctly by evaluating popular LLMs.
Contrary to previous work, we show that increasing the model size consistently improves the models’ ability to handle negations. Furthermore, we find that both the models’ reasoning accuracy and robustness to negation are language-dependent and that the length and explicitness of the premise have a greater impact on robustness than language.
Our datasets can facilitate further research and improvements of language model reasoning in multilingual settings.
Negation: A Pink Elephant in the Large Language Models’ Room?
Tereza Vrabcová  and Marek Kadlčík  and Petr Sojka  and Michal Štefánik  and Michal Spiegel Faculty of Informatics, Masaryk University
1 Introduction
The negation poses a fundamental problem for large language models, as shown by many studies Kassner and Schütze (2020); Truong et al. (2023a); García-Ferrero et al. (2023); Nadeem et al. (2024); Kalouli et al. (2022); Zhang et al. (2023). The phenomenon is called “NO syndrome” or “negation blindness”.
Vision-Language models do not understand negation either Alhamoud et al. (2025): when asked to provide a picture with no elephant in the room, they put an elephant there anyway. It appears that negation tokens (e.g., “not”) have a limited effect on the representations learned distributionally.
However, in tasks requiring logical reasoning, processing negation is critical for reliable performance.
Figure 1: We extend existing entailment datasets FEVER and SNLI with negated hypotheses in four languages (ENG, UKR, CZE, GER) and measure how negations degrade the accuracy of model reasoning.
In this paper, we evaluate LLMs on the text entailment (TE) task, exploring the ability of current LLMs to reason under negation. As in the known story about the elephant met by six blind monks, the problem of LLMs dealing with negation has been touched on from a number of different angles and points of view. Hermann et al. (2013) pioneeringly sees shortcomings in the ability of language models to capture logical operations such as negation in continuous vector space representation of compositional distributional semantics. Other points of view are the inclusion of more negative training data Helwe et al. (2022), modification of the input text itself to remove negation whilst keeping the overall meaning of the text Rezaei and Blanco (2024), or the fine-tuning of the instruction prompt to explicitly include the possible presence of negation in the texts Truong et al. (2023a). Ettinger (2020) assessed BERT’s Devlin et al. (2019) sensitivity to the meaning of negation by measuring the proportion of items in which the model assigns higher probabilities to true completions than to false ones. Kassner and Schütze (2020) studied negation and mispriming effects for pre-trained language models. Models memorize both positive and negative facts but fail to generalize to unseen facts. Dziri et al. (2023) investigate limits of LLMs across three representative compositional tasks.
In this paper, on the other hand, in the series of experiments, we pin down and separate the primary causes of the problem:
RQ1:
Can the scale of model size improve models’ ability to handle negations?
In one of the findings of Truong et al. (2023a), it is posited that the LLMs are unable to reason under negation, with the increased size of the LLM not being a mitigating factor. We review this claim by evaluating models from the same LLM family with different model sizes.
RQ2:
Can the adjustments in architecture and model refinements improve models’ capabilities to handle negations?
Furthermore, we evaluate whether the difference in architecture is a deciding factor in the model’s ability to understand negation.
RQ3:
Is the models’ capability to handle negations language-dependent?
As most studies concerning negation in LLMs focus on English datasets, we extend our scope to include multiple languages with differing levels of word-order flexibility.
2 Methodology
To study the effect of negation on the reasoning reliability of large language models, we evaluate them on textual entailment Putra et al. (2024). In this task, the model receives two claims, a premise and a hypothesis, and decides whether the hypothesis follows from the premise.
In our experiments, we inspect whether the presence of negation in the hypothesis causes degradation in the accuracy of the language model reasoning. For this purpose, we prepare a set of paired evaluation sentences that differ only in a negation of the primary verb. We call the variants hypotheses without negation and hypotheses with negation based on whether they contain a negation. Necessarily, if one follows from the premise, the other contradicts it, and vice versa. If the models’ reasoning is robust to negations, its accuracy should be the same between the two groups.
We base our evaluation on the FEVER Thorne et al. (2018a) and SNLI Bowman et al. (2015) datasets. From these English NLI datasets, we prepared our variants of two datasets in four languages through the pre-processing and translation processes. We have extended the original English datasets with negated versions of the hypotheses, using our negation pipeline. We have manually reviewed the correctness of these hypotheses, manually modifying the sentences in cases where the pipeline fell short. We then translated these English datasets using DeepL and validated the quality of translation by manually checking 100 random samples from each dataset. Furthermore, we also evaluated a preexisting Czech translation of the FEVER-NLI dataset, which has been fully verified by a human annotator; we denote this dataset in our results as CES FEVER MANUAL. More details of dataset preparation are described in Appendix A.
The evaluated models are LLama 3 Llama Team (2024b) with 3B, 8B, and 70B parameters Llama Team (2024a), Qwen 2.5 Yang et al. (2024) with 1.5B, 7B, and 72B parameters Qwen Team (2024), and Mistral Nemo 12B, Small 22B, and Large 123B Mistral AI (2024).
All of the models evaluated are publicly available state-of-the-arg models based on the transformer architecture Vaswani et al. (2017), have open weights, a multi-lingual pre-training, and are fine-tuned to follow user instructions. In the evaluation, we prompt the model to judge if a hypothesis follows from a premise or not. We let the model finish a response “The answer is:” with either True or False using a greedy decoding. To ensure the response follows the predefined structure, we employ Outlines Willard and Louf (2023) and LMFE Gat (2024). More details of the prompt format can be found in Appendix C.
3 Results
We summarize our results in this section. We find that accuracy and negation sensitivity improve with increasing model size and the information specificity of dataset premises, while variations in model architectures and languages are not statistically significant (for details, see Appendix D.
Figure 2: Robustness to negations across model sizes: Models’ accuracy difference on textual entailment when facing a hypothesis (i) including and (ii) not including a negation. The results are an aggregate across four languages (CES, DEU, ENG, UKR) measured for FEVER (solid line; left) and SNLI (dashed line; right) datasets.
Figure 3: Robustness to negations across languages: Accuracy difference of smaller ($\leq$7B; left) and large models ($\geq$70B; right) on textual entailment when facing a hypothesis (i) including and (ii) not including a negation. Inputs are identical across different languages. The results are aggregated across three models in each size group, measured for two datasets: FEVER (solid line) and SNLI (dashed line).
Figure 4: Robustness to negations in detail: Accuracy difference caused by negation for all 9 evaluated models across all 4 evaluated languages and our two new datasets. The bigger models consistently handle negation better.
Model size has a positive correlation with models’ ability to handle negation.
Figure 4 shows the accuracy change of different models for the hypotheses with and without negation. We found that all evaluated models perform better than a random guess and that the size of the model has a strong positive correlation with the model’s ability to correctly interpret negation. The larger models narrow the gap between the accuracies, with two outliers on the SNLI dataset: a slight increase on the middle-sized Llama model (8B) and a marked drop on the middle-sized Mistral model (22B).
The difference in our findings can be caused by a particular choice of the evaluated models. However, a plausible explanation is that measured accuracy strongly benefits from constraining output to a set of predefined options rather than expecting the language model to adhere by itself to a desired output format.
The overall findings are in contrast to the Finding of Truong et al. (2023b), where it is posited that pre-trained language models perform comparably or worse than a random baseline, and model scaling has almost no effect.
Larger LMs are more sensitive to negation for TE task.
As we can see in Table 5 in the Appendix, the size of the model has a positive correlation with the overall negation sensitivity of the models. Larger models are less likely to classify pair hypotheses with the same textual entailment label, and the impact of negation is more effectively propagated through the model. One outlier in our results is the Mistral (22B) model, which has outperformed the Mistral (12B) model on CES, ENG, and UKR SNLI datasets.
Architecture adjustments do not substantially impact models’ ability to handle negation.
Apart from differences in size, performance across different LLM families are comparable.
This suggests that, on average, current architectural differences in state-of-the-art models may not be the deciding factor in a model’s ability to understand negation. Instead, it seems that the size of the model plays a prominent role in determining how well it handles negations, likely due to its increased capacity to capture nuanced language patterns. In Appendix D.4, we further discuss the influence of model size on the models’ ability to more accurately model the language patterns.
Models’ accuracy is not language-dependent.
Figure 4 shows the accuracy change on different models with regard to the language of the dataset. Our results do not fully conform to our initial presuppositions that the models will perform worse in languages with higher word-order flexibility. The Czech datasets have generally performed worst out of the evaluated datasets, with the CES FEVER MANUAL achieving higher accuracy on hypotheses without negation than the CES FEVER dataset.
However, in conjunction with Figures 4 and Figure 4, our results lead us to conclude that the length and specificity of the premise have a larger impact than the language of the evaluated dataset. Whereas the models were able to properly handle negation on the FEVER dataset, which is based on a fact verification dataset and thus has rich and factually specific premises, the negation on the SNLI dataset, which is more dependent on common sense knowledge and implicit semantical patterns the models learn during training, has shown itself to be problematic, even for large models.
4 Conclusions and future work
This work presents a novel dataset enabling the evaluation of language models’ robustness in handling negations in four languages.We have extended an existing FEVER Thorne et al. (2018b) and SNLI Bowman et al. (2015) datasets to include paired hypotheses that are logically opposite to each other by negating the primary verb in the original hypothesis. We have evaluated the models’ accuracy separately on hypotheses without negation and hypotheses with negation computed the accuracy drop that occurred due to the addition of negation in the hypothesis.
We find that model size plays a key role in handling negation correctly, countering the conclusions of previous work Truong et al. (2023a). However, there seems to be minimal effect on the accuracy that can be attributed to the different architectures and pre-training data. The results on differing language datasets only partially match the initial expectations, with models performing worst on the Czech configuration; however, datasets created using machine translation performed similarly, regardless of the language morphology. The accuracy more heavily depended on the specificity of information in the premise; the more common sense SNLI dataset performed worse than the more fact-explicit FEVER dataset.
Our work provides new resources for addressing the deficiency of language models in handling negations. Our methodology will be used to find the root cause of the negation problem by computing the accuracy on subsets of our new datasets in 4 languages (e.g., a subset of nonprojective expression of negation, a subset of different ways of expressing syntactic or semantic negation Kalouli et al. (2022) or to evaluate sub-clausal negation based on a dataset of Truong et al. (2022). We will use the lense of LLM-microscope Razzhigaev et al. (2025) to look on the tokens that play decisive role.
Limitations
The details of model preparation (training data, hyperparameters, fine-tuning) limit the full understanding of the results we report. Also, the results might be biased by the [semi]automated translations of datasets to their German, Czech, and Ukrainian versions. However, we did a manual check of the quality of a subset of the translations and did not encounter any errors related to the application of negation in context.
Appendix A Dataset preparations
We evaluate the accuracy of the models in our experiments using two datasets formatted for the NLI task, i. e. containing premise-hypothesis pairs and the label denoting their logical relationship (entailment, neutral, contradiction):
SNLI
Bowman et al. (2015) is a popular NLI dataset focusing on the evaluation of general and common sense knowledge of the LLMs. The premises are captions of images capturing everyday situations and, as such, are quite short, consisting of one to two sentences. The hypotheses have been generated manually by human annotators.
FEVER-NLI
Thorne et al. (2018a) is a modification of the FEVER dataset, which originally focused on the task of fact extraction and verification. The dataset consists primarily of factual statements retrieved from Wikipedia.
As our aim is to study the effects of negation on the accuracy of the models, we extend these datasets to contain the negation of the original hypothesis. The original datasets contain a low percentage of hypotheses with common negation markers, such as no, n’t, not, nobody, neither, or nothing, and differ in the verbosity of premises, as seen in the Table 1.
In Table 2, we include average lengths of premises and hypotheses of the newly prepared datasets for comparison.
SNLI
FEVER-NLI
Rows in the dataset 20,000 10,000
Rows w/ negation 868 189
% of rows w/ negation 4.34 1.89
Premise word count 13.94 51.96
Hypothesis word count 07.50 08.80
Table 1: Comparison of the original English SNLI and FEVER-NLI datasets
CES
FEVER
MANUAL
CES
FEVER
CES
SNLI
ENG
FEVER
ENG
SNLI
DEU
FEVER
DEU
SNLI
UKR
FEVER
UKR
SNLI
Premise (chars) 268.36 265.06 51.71 266.31 58.02 306.60 70.33 278.54 58.55
Hypothesis without negation (chars) 42.87 41.69 24.06 42.12 28.71 49.63 34.55 43.38 27.71
Hypothesis with negation (chars) 44.89 43.58 26.42 51.13 37.26 53.69 38.87 45.11 29.70
Premise (words) 49.36 48.85 10.74 54.12 14.05 55.43 13.58 47.90 10.55
Hypothesis without negation (words) 7.86 7.66 5.20 8.72 7.13 8.97 7.01 7.35 4.99
Hypothesis with negation (words) 7.86 7.69 5.32 10.95 9.23 9.82 7.93 8.34 5.99
Table 2: Comparison of average lengths across evaluated datasets
A.1 Preparation of new datasets in English
In order to prepare new datasets containing negation, we first sanitize the datasets, and then we extend the datasets with hypotheses, including negation. We then manually replace incorrectly generated sentences and create two datasets: the EngNoFEVER dataset, containing 2,600 premise-hypothesis-negated_hypothesis triplets, and the EngNoSNLI dataset, containing 6,261 triplets.
Sanitization and Filtration
Firstly, we sanitize the datasets by removing all incomplete or invalid rows, i.e., removing rows with empty premises, empty hypotheses, or invalid labels. Secondly, we remove rows that have the neutral entailment label as the effect of negation on their entailment value is not uniform.
Negation
We employ the negate Python module for our negation. This module is a rule-based negation tool for the English language, utilizing the spaCy model en_core_web_md for POS tagging and dependency parsing in order to correctly locate the verb in the sentence. The verb is then negated by the inclusion of an auxiliary verb and a negation marker, with our preference in the module set to the contraction n’t. The tool is able to negate in both directions, meaning that it can remove negation from the verb as well.
Rows with hypotheses that do not contain verbs are discarded during the generation process.
Manual verification
As the tool we used for the hypothesis generation is rule-based, it comes with certain limitations and cases that are not covered by the rules and have to be fixed manually.
Some, any, nor yet are properly supported.
Example 1:
Input: There is someone jumping.
Negate output: There isn’t someone jumping.
Manual change: There isn’t anyone jumping.
Example 2:
Input: Keegan-Michael Key has yet to play anyone.
Negate output: Keegan-Michael Key doesn’t have yet to play anyone.
Manual change: Keegan-Michale Key has played someone.
The rules of the module as aimed the Not-negation, most likely due to its higher frequency in the English language, in comparison to the No-negation.
Example 3:
Input: There is no one running.
Negate output: There isn’t no one running.
Manual change: There is someone running.
The typos in the original datasets have been fixed only in cases where the typo directly impacted the negation of the verb.
A.2 Translation of new datasets into Czech, German, and Ukrainian
To explore the impact of negation on the accuracy of the LLMs across languages, we translate our English (ENG) datasets into Czech (CES), German (DEU), and Ukrainian (UKR). We have chosen these languages due to their differing features, such as morphology, orthography, alphabets, and projectivity, as specified in Table 3.
ENG
CES
DEU
UKR
Language Family West Germanic West Slavic West Germanic East Slavic
Script Latin Latin Latin Cyrillic
Linguistic Typology Analytic Fusional Fusional w/ Agglutinative Features Fusional
Word Order Flexibility Low High Moderate High
Table 3: Comparison of linguistic features between languages English (ENG), Czech (CES), German (DEU), and Ukrainian (UKR).
Translator Selection
There are many machine translation tools to choose from. To choose between Google Translate and DeepL, we have translated 100 rows from the modified English FEVER-NLI dataset into Czech, comparing the quality of translation between the translations and to a preexisting translation of the FEVER-NLI dataset that has been manually verified by a human annotator. From these 100 rows, DeepL provided higher quality for 44 rows, Google Translate for 30, with 26 rows having equal translation quality between them.
As the difference between the translations is not statistically significant in the 100 examples, we have chosen to use DeepL to translate the datasets due to the higher percentage of winning translations.
Translation validation
After we had translated the datasets into our target languages, we randomly selected 100 rows from each dataset and checked the quality of the translation.
Although the translation quality of the model we have used is generally of high quality, it is still an LLM that is not immune to hallucination on rarely seen data. We have, for example, in the Ukrainian SNLI dataset, observed the outlier case shown in Table 4. This behavior does not occur when the sentence does not include negation. We suppose that the combination of the inclusion of negation, the tense in the original English sentence, and the inclusion of an initialism (pb and j) created confusion in the model. However, as the training data of the model isn’t public, we cannot confirm our hypothesis.
Original/Translated Sentence
Sentence in English
ENG A man isn’t eating pb and j A man isn’t eating pb and j
CES Muž nejí pb a j. A man isn’t eating pb and j.
DEU Ein Mann isst kein Pb und J A man isn’t eating Pb and J
UKR Чоловiк не їсть пломбiр з соєвими бобами A man isn’t eating ice cream with soy beans
Table 4: Translation of an outlier example sentence using DeepL
Appendix B Full details of accuracy across evaluated models
Figure 5: Absolute accuracy for all evaluated models across all evaluated languages and datasets for textual entailment of hypotheses without negation.
Figure 6: Absolute accuracy for all evaluated models across all evaluated languages and datasets for textual entailment of hypotheses with negation.
We include the full tables of accuracies of models for hypotheses with and without negation on the task of textual entailment. In both tables, we can see that almost all models achieve the highest accuracy on the English datasets, on both hypotheses, with and without negation.
However, the accuracy of the models from the Qwen family shown in Figure 4 gave us a pause, particularly the fact that overperformance is present only for one specific dataset. The English SNLI dataset with hypotheses without negation is very similar to the original SNLI dataset, as less than 5 percent of hypotheses in the original test dataset contain negation. As the original SNLI is a very popular benchmark for the evaluation of NLI and textual entailment tasks, and the Qwen overperformance did not translate across languages on the same dataset, nor did it translate into an increase of accuracy on hypotheses with negation, we have to wonder whether the training data of the Qwen model family inadvertently included the SNLI evaluation data.
Model (size)
CES
FEVER
MANUAL
CES
FEVER
CES
SNLI
ENG
FEVER
ENG
SNLI
DEU
FEVER
DEU
SNLI
UKR
FEVER
UKR
SNLI
Llama (3B) 68.04 % 68.04 % 64.88 % 73.54 % 81.23 % 49.08 % 61.85 % 65.19 % 68.36 %
Llama (8B) 80.5 % 80.5 % 75.87 % 88.62 % 86.93 % 78.07 % 84.46 % 80.38 % 79.08 %
Llama (70B) 87.62 % 87.62 % 80.95 % 94.81 % 92.29 % 84.22 % 89.69 % 84.77 % 78.98 %
Qwen (1.5B) 55.92 % 55.92 % 40.44 % 64.77 % 62.37 % 48.08 % 58.58 % 54.31 % 37.85 %
Qwen (7B) 81.85 % 81.85 % 71.03 % 89.54 % 77.35 % 70.95 % 83.85 % 79.46 % 72.85 %
Qwen (72B) 87.65 % 87.65 % 71.83 % 92.65 % 80.93 % 75.56 % 88.38 % 85.38 % 70.34 %
Mistral (12B) 80.65 % 80.65 % 69.4 % 87.88 % 87.06 % 74.8 % 83.81 % 77.04 % 66.19 %
Mistral (22B) 85.23 % 85.23 % 62.11 % 91.46 % 66.79 % 63.79 % 87.08 % 77.81 % 62.08 %
Mistral (123B) 87.81 % 87.81 % 72.1 % 91.31 % 80.59 % 77.3 % 89.08 % 84.23 % 74.09 %
Table 5: Evaluation of the negation sensitivity of each model across individual datasets. Our negation sensitivity metric represents the percentage of datasets in which the model assigned opposite entailment scores to pairs of hypotheses with and without negation. The best results (in bold) were achieved only by the biggest models supporting our RQ1.
Appendix C Prompting template
All the models we evaluated support a chat-like interface with a system prompt, a user prompt, and a model response.
As a system prompt, we used the following template:
You are a fact checker for queries in the <English / Czech / German / Ukrainian> language. You will be given a premise, which you know is factually correct, and a hypothesis. You will return the truth value of the hypothesis, based on the premise. Return True if the hypothesis is correct and False if the hypothesis is incorrect.
The user prompt follows a simple format:
Premise: <premise>
Hypothesis: <hypothesis>
The model decides by finishing a prepared response:
The answer is: <True / False>
Appendix D Statistical tests
In this section, we include the computation of the statistical tests upon which we base our claims in Section 3. Unless otherwise stated, for our computations, we are using alpha level (significance threshold) $\alpha=0.05$.
D.1 Model size and relative accuracy change
Using the results shown in Figure 4, we compute the correlation of the model size and the relative accuracy change per each evaluated dataset and overall. Using the Shapiro-Wilk test, we find that the majority of datasets do not have normal distribution over the different model sizes. Thus, we compute correlation using the Spearman correlation coefficient, as the Pearson correlation coefficient presumes a normal distribution of values. Values used during computation are shown in Table 6.
The results indicate a strong positive correlation between model size and the relative accuracy change on the FEVER dataset and a weak-to-moderate positive correlation on the SNLI dataset. For an averaged relative accuracy change for all languages per model, the results indicate a strong positive correlation with the Spearman correlation coefficient of $0.867$.
We posit that the differences in correlation are due to different philosophies behind the datasets. The FEVER dataset was originally based on the fact verification dataset and consists of factually specific premises that are sufficient for the prediction of the textual entailment. The SNLI dataset, on the other hand, relies more heavily on implicit semantical patterns that the models learn during training, introducing an uncertainty we cannot easily predict.
CES
FEVER
MANUAL
CES
FEVER
CES
SNLI
ENG
FEVER
ENG
SNLI
DEU
FEVER
DEU
SNLI
UKR
SNLI
UKR
FEVER
Shapiro-Wilk test $p$-value 0.279 0.004 0.008 0.134 0.001 0.016 0.046 0.003 0.016
Spearman correlation coefficient 0.867 0.883 0.400 0.883 0.250 0.900 0.300 0.883 0.150
Table 6: Values used during the computation of correlation between the model size and the relative accuracy change across all evaluated datasets.
D.1.1 Negation sensitivity
Similarly to the relative accuracy change, we compute the correlation between model size and the model’s sensitivity to negation. As per the computed $p$-values for the Shapiro-Wilk test, we can see that the majority of our negation sensitivity metric across datasets does not have a normal distribution either. Therefore, we are again using the Spearman correlation.
The results indicate a very strong positive correlation of negation sensitivity with the model size on the FEVER datasets, with a rather weak correlation on the SNLI datasets. For an average negation sensitivity for all languages per model, the results indicate a strong positive correlation, with a Spearman correlation coefficient of $0.866$.
CES
FEVER
MANUAL
CES
FEVER
CES
SNLI
ENG
FEVER
ENG
SNLI
DEU
FEVER
DEU
SNLI
UKR
SNLI
UKR
FEVER
Shapiro-Wilk $p$-value 0.013 0.013 0.052 0.008 0.405 0.101 0.002 0.027 0.022
Spearman correlation w/ model size 0.950 0.950 0.583 0.817 0.267 0.683 0.900 0.817 0.400
Table 7: Values used during the computation of correlation between the model size and the negation sensitivity across all evaluated datasets.
D.2 Model family
As the model family is a categorical variable, instead of correlation, we compute and compare the variance of the relative accuracy change across the three model families (Qwen, Mistral, and Llama), using values shown in Figure 4 and the Kruskal-Wallis test as our analysis method. Our null hypothesis is that all model families have the same average of relative accuracy changes, and the alternative hypothesis is that the averages are not the same.
[…]
D.3 Language
We compute and compare the variance of the relative accuracy change across the four evaluated languages (CES, ENG, DEU, UKR), similarly to the model family variable. We aggregate values shown in Figure 4 per language and use the Kruskal-Wallis test as our analysis method. Our null hypothesis is that all languages have the same average of relative accuracy change, and the alternate hypothesis is that the averages are not the same.
[…]
As the $p$-value is not less than our alpha level, we cannot reject our null hypothesis. Thus, we conclude that there is no significant difference between the relative accuracy change across different languages.
D.4 Sentence complexity
As mentioned in Section 3, we posit that the larger models can more accurately capture nuanced language patterns, enabling them to correctly handle negations in more complex sentences that elude the smaller models. From each model family, we take the smallest and the largest models and compare the average sentence complexity of the hypotheses the model is able to correctly classify. However, we hypothesize that many hypotheses correctly classified by the smaller model are also correctly classified by the larger model. Therefore, for the larger model, we compute the average sentence complexity of the additional correctly classified hypotheses, i.e., the subset for the larger model constitutes of hypotheses the larger model correctly classified which are not present in the smaller model’s subset.
We evaluate the sentence complexity using the following metrics:
Hypothesis depth.
Depth of the hypothesis dependency tree. Sentence complexity increases with the increase of the depth of the dependency tree.
Tables 8 and 9 show the the increase of average hypothesis depth between the smallest and the largest model per the language model family. Due to the relative simplicity of the hypotheses (short length as described in Table 2, average of one verb per hypothesis), the increase of depth is not substantial, being less than one level across all datasets for all language model families.
Dissimilarity of hypothesis to the premise.
Number of content words present only in the hypothesis and not the premise, divided by the total number of content words in the hypothesis. Sentence complexity increases with the increase of dissimilarity. We compute the similarity on the lemmatized forms of the words with the most common stop words per each language filtered out. High dissimilarity may be caused by the usage of synonyms, hypernyms, hyponyms, or other linguistic features that require higher understanding of syntactic patterns in the hypothesis.
Tables 10 and 11 show the increase in the average dissimilarity of the premise and its hypothesis between the smallest and the largest model per the language model family. Here we can see a marked increase in dissimilarity for hypotheses with negation, especially within the Qwen family models, showcasing the largest model’s increased ability in correctly reason and determine entailment on texts that use different formulations to express the same meaning.
CES
FEVER
MANUAL
CES
FEVER
CES
SNLI
ENG
FEVER
ENG
SNLI
DEU
FEVER
DEU
SNLI
UKR
FEVER
UKR
SNLI
Llama Family 0.027 0.215 0.064 0.270 0.196 0.219 0.056 0.155 0.049
Qwen Family 0.082 0.173 0.006 0.215 0.174 0.024 0.030 0.148 0.041
Mistral Family -0.016 0.184 0.037 0.271 0.137 0.245 0.001 0.121 -0.017
Table 8: Increase of average depth of hypotheses without negation that were correctly predicted only by the largest model of the LM family and the correctly predicted hypotheses by the smallest model of the LM family.
CES
FEVER
MANUAL
CES
FEVER
CES
SNLI
ENG
FEVER
ENG
SNLI
DEU
FEVER
DEU
SNLI
UKR
FEVER
UKR
SNLI
Llama Family 0.019 0.076 0.090 0.089 0.145 0.041 0.030 0.150 0.067
Qwen Family 0.038 -0.09 0.027 -0.016 0.065 -0.053 -0.084 -0.006 -0.080
Mistral Family -0.016 0.031 0.064 0.125 0.101 0.103 -0.012 0.151 -0.017
Table 9: Increase of average depth of hypotheses with negation that were correctly predicted only by the largest model of the LM family and the correctly predicted hypotheses by the smallest model of the LM family.
CES
FEVER
MANUAL
CES
FEVER
CES
SNLI
ENG
FEVER
ENG
SNLI
DEU
FEVER
DEU
SNLI
UKR
FEVER
UKR
SNLI
Llama Family 2.52 % -0.12 % -4.31 % -5.77 % -15.30 % -4.73 % -16.01 % -1.20 % -6.83 %
Qwen Family -0.47 % 4.19 % 2.55 % 7.88 % -5.00 % 0.90 % 3.58 % 4.66 % 1.44 %
Mistral Family 0.89 % 4.49 % 0.42 % 6.94 % -1.63 % 1.02 % 6.26 % 2.49 % -1.92 %
Table 10: Increase of average sentence dissimilarity of premises and the evaluted hypotheses without negation that were correctly predicted only by the largest model of the LM family and the correctly predicted hypotheses by the smallest model of the LM family.
CES
FEVER
MANUAL
CES
FEVER
CES
SNLI
ENG
FEVER
ENG
SNLI
DEU
FEVER
DEU
SNLI
UKR
FEVER
UKR
SNLI
Llama Family 2.00 % 3.97 % -3.02 % 2.36 % 0.62 % 5.52 % 21.12 % 5.56 % -1.52 %
Qwen Family 0.24 % 12.49 % 19.41 % 12.50 % 21.14 % 10.56 % 25.16 % 9.44 % 15.28 %
Mistral Family 1.59 % 4.06 % 11.77% 5.22 % 6.11 % 5.30 % 11.96 % 4.94 % 10.36 %
Table 11: Increase of average sentence dissimilarity of premises and the evaluated hypotheses with negation that were correctly predicted only by the largest model of the LM family and the correctly predicted hypotheses by the smallest model of the LM family.